# Collective Variable Free Transition Path Sampling with Generative Flow Network

### v2

@@ 3.2 TPS-GFN: GFlowNets for transition path sampling
In this section, we explain our approach, called TPS-GFN, that amortizes the expensive cost of sampling transition paths from the distribution $p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})$. Our key idea is threefold: (1) using the trajectory balance (TB) objective to train a generator $p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0})=\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t}|\bm{x}_{t-1})$ towards the true distribution $p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})$, (2) relaxation of the target distribution, and (3) parameterization of the generator with the bias potential. We provide a full description of the algorithm in Algorithm 1.
@@
for all trajectories $\bm{x}_{1:T}$. To this end, we propose to use the one-step TB objective [17], the squared log-ratio between the unnormalized generator $Z_{\theta}p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0})$ and the unnormalized target distribution $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})\prod_{t=1}^{T}p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})$, as follows:
$\mathcal{L}_{\text{TB}}(\bm{x}_{1:T};\theta)=\left(\log\frac{Z_{\theta}\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t}|\bm{x}_{t-1})}{\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})\prod_{t=1}^{T}p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})}\right)^{2},$ (2)
where $Z_{\theta}\in\mathbb{R}$ is a learnable scalar parameter to estimate the normalization constant $Z$. We match the generator with the target distribution by minimizing the loss function $\mathcal{L}_{\text{TB}}(\bm{x}_{1:T};\theta)$ over trajectories $\bm{x}_{1:T}$ sampled from an arbitrary training distribution, i.e., off-policy training. We provide an explicit derivation of the one-step TB objective (Equation 2) in Appendix A.
To leverage the ability of the TB objective, we use a replay buffer that collects the past samples from the tempered version of the generator and reuses them for training. This improve sample efficiency and stability of training, both for GFlowNets [15] and reinforcement learning [40]. TPS-GFN alternates between (1) sampling from the generator to store molecular trajectories in the replay buffer and (2) training the generator on samples from the replay buffer, as described in Algorithm 1.
@@
1:  Initialize an empty replay buffer $\mathcal{B}$, the biased potential $b_{\theta}(\bm{r})$, and the parameter $Z_{\theta}$.
@@
3:     Sample $M_{1}$ trajectories $\{\bm{x}_{0:T}^{(m)}\}^{M_{1}}_{m=1}$ from the tempered version of biased MD (Equation 3).
@@
6:        Sample $M_{2}$ paths $\{\bm{x}_{0:T}^{(m)}\}^{M_{2}}_{m=1}$ from the replay buffer $\mathcal{B}$.
7:        Update parameters $\theta$ to minimize the loss $\mathcal{L}$ (Equation 2) using the samples $\{\bm{x}_{0:T}^{(m)}\}^{M_{2}}_{m=1}$.
@@
As shown by Malkin et al. [37], the TB objective automatically performs gradient variance reduction for variational inference by estimating a marginal quantity (the partition function) that acts as a baseline and allows off-policy learning without the need for reweighted importance sampling. This is also a slight modification to the VarGrad objective [41, 42] that was proposed to reduce the gradient of evidence lower bound (ELBO).
We note the similarity of our TPS-GFN with the prior work [14], which minimizes the Kullback-Leibler (KL) divergence between the generator and the target distribution. In comparison, the TB objective yields better training dynamics due to reduced variance in the gradients, and better sample efficiency due to the ability to reuse the collected sample with an external replay buffer. Furthermore, TPS-GFN does not suffer from the mode-seeking behavior of KL divergence.
Target distribution relaxation. Learning to imitate the target distribution $p_{\mathcal{A},\mathcal{B}}$ with TB objective is challenging since its probability is rarely non-zero. To alleviate this issue, we relax the target distribution by (1) allowing trajectories to reach the meta-stable region $\mathcal{B}$ before termination and (2) relaxing the binary indicator function into a smooth function. In particular, we relax the target distribution into the following form:
@@
where $\tilde{Z}$ is the normalizing constant and $\tilde{\mathbf{1}}_{\mathcal{B}}(\bm{r}_{1:T})$ is the relaxed indicator function defined as follows:
@@
$\tilde{\mathbf{1}}_{\mathcal{B}}(\bm{r}_{1:T})=\max_{t\in\{1,\ldots,T\}}\exp\left(-\frac{\lVert D(\bm{r}_{t})-D(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}}{2\sigma^{2}}\right).$
Here, $\lVert\cdot\rVert_{F}$ is the Frobeneous norm, $\sigma>0$ is a hyperparameter to control the degree of relaxation, and $D(\bm{r})\in\mathbb{R}^{N\times N}$ is a distance matrix with $(i,j)$-th element denoting the Euclidean distance $\lVert r_{i}-r_{j}\rVert_{2}$ between the $i$-th and the $j$-th atoms [43].
Parameterizing generator with bias potential. Similar to Holdijk et al. [14], we model the generator $p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0})$ as the discretization of the following biased Langevin dynamics:
$\mathrm{d}\bm{r}=\bm{v}\mathrm{d}t,\quad\mathrm{d}\bm{v}=-\frac{\nabla_{\bm{r}}(U(\bm{r})+b_{\theta}(\bm{r}))}{\bm{m}}\mathrm{d}t-\gamma\bm{v}\mathrm{d}t+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\mathrm{d}\bm{w},$
where $b_{\theta}(\bm{r})$ is the bias potential that guides the Langevin dynamics towards the meta-stable region $\mathcal{B}$. We discretize the biased Langevin dynamics as follows:
$\bm{r}_{t+1}=\bm{r}_{t}+\bm{v}_{t}\Delta,\qquad\bm{v}_{t+1}=(1-\gamma\Delta)\bm{v}_{t}+\frac{-\nabla_{\bm{r}}(U(\bm{r}_{t})+b_{\theta}(\bm{r}_{t}))}{\bm{m}}\Delta+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\bm{\varepsilon}_{t}\Delta,$ (3)
We model the bias potential $b_{\theta}(\bm{r})$ or directly bias force $\mathbf{b}_{\theta}(\bm{r})=\nabla_{\bm{r}}b_{\theta}(\bm{r})$ for the generator. We use a multi-layered perception (MLP) architecture to parameterize the bias potential or force.
4 Experiment
Table 1: Benchmark scores on Alanine Dipeptide which have well-known CVs. All metrics are averaged over 64 trajectories. ETP and EFP are reported only if the final state is in the target meta-stable state. The best values are highlighted in bold except for MD. Values from Holdijk et al. [14] are indicated by $\dagger$. Methods using bias force and potential are denoted by (F) and (P), respectively.
[…]
MD (300K) $7.78\pm 0.61$ $0.00$--- 0--5.02 $\pm$ 0.15 088.27±5.65plus-or-minus88.275.6588.27$\pm$ 5.6588.27 ± 5.650 0.--
MD (9000K) $7.26\pm 2.47$ 03.12 --770.18 $\pm$ 151.85 $-25.47\pm 4.91$ $101.82\pm 31.02$ $\phantom{-}604.02\pm 136.73$
PIPS (F) $2.06\pm 0.27$ $56.25$-000.48 $\pm$ 10.310 0--4.49 $\pm$ 0.36 012.15±3.49plus-or-minus12.153.4912.15$\pm$ 3.4912.15 ± 3.490 $-15.85\pm 11.21$
PIPS (P) $1.21\pm 0.31$ $63.50$ 00--8.35 $\pm$ 8.0400 00.-- 0-- 0.--
[…]
TPS-GFN (P) 1.02 $\pm$ 0.08 98.43 0--16.62 $\pm$ 5.1000 0--$4.85\pm 0.12$ 004.91 $\pm$ 1.700 045.18±6.88plus-or-minus45.186.88-45.18$\pm$ 6.88- 45.18 ± 6.8800
Table 2: Benchmark scores on Polyproline and Chignolin. All metrics are averaged over 64 trajectories. The best values are highlighted in bold except for MD. $\ddagger$ indicates reproduced results from the prior work [14]. Methods using bias force and potential are denoted by (F) and (P), respectively.
Molecule Method EPD ($\downarrow$) LL ($\uparrow$) EPCD ($\downarrow$)
$\mathrm{nm\times 10^{-3}}$$\times 10^{2}$ $\mathrm{nm}$
MD (300K) $18.83\pm 1.59$-00 4.45 $\pm$ 0.57 0 $82.43\pm 3.69$
PIPS (F) $29.13\pm 2.94$ $-215.82\pm 87.03$ $111.46\pm 4.86$
Polyproline PIPS (P) $12.21\pm 1.77$ $-210.27\pm 96.11$ $76.90\pm 6.17$
TPS-GFN (F) 07.94 $\pm$ 2.13 006.93±0.98plus-or-minus6.930.98-6.93$\pm$ 0.98- 6.93 ± 0.980 42.85 $\pm$ 8.00
TPS-GFN (P) $13.86\pm 1.08$-002.64 $\pm$ 0.520 $83.23\pm 4.59$
MD (300K) $381.68\pm 48.74$-010.11 $\pm$ 1.030 $259.63\pm 12.58$
PIPS (F) 068.02 $\pm$ 20.63 $-251.19\pm 63.00$ $188.56\pm 29.77$
Chignolin PIPS (P) 068.87 $\pm$ 9.390 $-241.22\pm 58.15$ $186.30\pm 11.79$
[…]
TPS-GFN (P) 026.31 $\pm$ 1.810 00--0.88 $\pm$ 1.420 101.43 $\pm$ 7.270
4.1 Molecular systems
In this section, we evaluate our TPS-GFN on three molecular systems, (1) Alanine Dipeptide, (2) Polyproline Helix, and (3) Chignolin. We run molecular dynamics (MD) simulations on a femtosecond ($\mathrm{fs}$) scale to sample paths and evaluate them at the temperature of 300K, using the OpenMM library. We report more details about OpenMM and model configurations in Section B.1 and Section B.2, respectively.
Alanine Dipeptide. The Alanine Dipeptide is a small peptide consisting of two alanine residues linked by a peptide bond. Alanine Dipeptide has a well-known collective variable (CV), i.e., two dihedral angles $(\phi,\psi)$. We sample the transition paths from the meta-stable region $C5$ to $C7_{ax}$ for a 500 $\mathrm{fs}$.
Polyproline Helix and Chignolin. In our experiments, the Polyproline Helix consists of 3 proline residues while the Chignolin consists of 10 amino acid residues. Their CVs and mechanisms are less understood than Alanine Dipeptide due to their complexity. We sample transition paths of Polyporline between the meta-stable region PP-II and PP-I for a 10000 $\mathrm{fs}$, and Chignolin folding process for a 5000 $\mathrm{fs}$. In training, we use the tempered version of biased MD simulations at a temperature of 600K.
(a) Molecular dynamics (MD)
@@ 4.1 Molecular systems
Figure 3: Visualization of 64 paths starting in meta-stable region $C5$ on the Ramachandran plot of Alanine Dipeptide. White circles indicate meta-stable states, and stars indicate transition states. (3(a)) The paths from unbiased MD simulations that fail to escape the initial meta-stable region. (3(b)) The paths generated by PIPS, passing through only one transition state. (3(c)) The paths generated by TPS-GFN, passing through both transition states. For clarity, 10 paths are highlighted.
@@ 4.2 Quantitative results
In our experiments, we compare our method with unbiased MD and PIPS [14]. To compare the generated transition paths, we consider three metrics following the prior work [14]. Expected pairwise distance (EPD) measures the similarity between the final conformation and the target conformation by comparing their interatomic distances. Similar to EPD, target hit percentage (THP) evaluates how many paths achieve the target meta-stable region in a binary manner. Unlike EPD and THP which only focus on the final conformation of the transition path, energy transition point (ETP) measures the maximum potential energy of a transition path, referring to the ability of a model to find the transition states when crossing the energy barrier.
@@
However, ETP only evaluates the potential energy of a transition state, it can not reflect the realisticity of an overall transition path. To resolve this issue, we propose log-likelihood (LL) which refers to log-likelihood from the unbiased MD simulation. Similarly, since EPD and THP only consider the distance between the final conformation of sampled paths and the target conformation, they can not reflect the potential energy of the final conformation which explains its stability and realisticity. To evaluate it, we propose the potential energy of the final point (EFP), and the expected pairwise Coulomb distance (EPCD) which consider both the potential energy of the final conformation and the distance between the final conformation of sampled paths and the target conformation. EPCD measures both geometric and electrostatic similarity between the final conformation and the target conformation using the Coulomb matrix. The Coulomb matrix captures atomic mass and energy level around the final state better than the distance matrix. More details of metrics are in Section B.3.
@@
Alanine Dipeptide. Since Alanine Dipeptide has well-known CVs, i.e., two dihedral angles $(\phi,\psi)$, we can evaluate THP, ETP, and EFP which requires defining the meta-stable region with a CV. In Table 1, our TPS-GFN outperforms the baselines for both previous metrics (EPD, THP, ETP) and proposed metrics (LL, EPCD, EFP) regardless of predicting bias potential or force. This shows that the TB objective used in our method leads to more realistic and high-quality transition paths than KL divergence used in PIPS. In particular, the $1.02$ EPD and $98.43\%$ THP score of TPS-GFN show that our method faithfully amortizes the rejection sampling from unbiased MD simulations. We also highlight that the EFP of our method is much closer to the potential energy of the target conformation $\bm{r}_{C7_{ax}}$, $-84.86$ $\mathrm{kJmol^{-1}}$, than the baselines.
@@
Polyproline Helix and Chignolin. Unlike Alanine Dipeptide, the CVs of Polyproline Helix, and Chignolin are less understood. Therefore, we exclude metrics THP, ETP, and EFP which require CVs to define the meta-stable state. In Table 2, our method consistently outperforms baselines in both Polyproline Helix and Chignolin, regardless of predicting bias potential or force. Our model samples more realistic transition paths than the baselines with respect to the overall path and final conformation. We note that as the size of the molecule is larger, the EPD and EPCD performances decreases. For the large size of molecules including Chignolin, the bias potential produces better transition paths than bias force since the output of bias potential is always scalar regardless of the size of the molecule and the external force from the bias potential is conservative.

### v3

@@ 3.2 TPS-GFN: GFlowNets for transition path sampling
In this section, we explain our approach, called TPS-GFN, that amortizes the expensive cost of sampling transition paths from the distribution $p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})$. Our key idea is threefold: (1) using the trajectory balance (TB) objective to train a generator $p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0})=\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t}|\bm{x}_{t-1})$ towards the true distribution $p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})$, (2) relaxation of the target distribution, and (3) parameterization of the generator with the neural bias potential. We provide a full description of the algorithm in Algorithm 1.
@@
for all paths $\bm{x}_{1:T}$. To this end, we propose to use the one-step TB objective (Malkin et al., 2022a), the squared log-ratio between the unnormalized generator $Z_{\theta}p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0})$ and the unnormalized target distribution $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})\prod_{t=1}^{T}p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})$, as follows:
$\mathcal{L}_{\text{TB}}(\bm{x}_{1:T})=\left(\log\frac{Z_{\theta}\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t}|\bm{x}_{t-1})}{\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})\prod_{t=1}^{T}p_{\text{MD}}(\bm{x}_{t}|\bm{x}_{t-1})}\right)^{2},$ (2)
where $Z_{\theta}\in\mathbb{R}$ is a learnable scalar parameter to estimate the normalization constant $Z$. We match the generator with the target distribution by minimizing the loss function $\mathcal{L}_{\text{TB}}(\bm{x}_{1:T})$ over paths $\bm{x}_{1:T}$ sampled from an arbitrary training distribution, i.e., off-policy training. We provide an explicit derivation of the one-step TB objective (Equation 2) in Appendix A.
@@
1:  Initialize an empty replay buffer $\mathcal{B}$, a neural biased potential $b_{\theta}$, a scalar parameter $Z_{\theta}$, and a temperature schedule $\lambda_{\text{start}}=\lambda_{1}>\cdots>\lambda_{I}=\lambda_{\text{end}}$.
@@
3:     Generate $M_{1}$ paths $\{\bm{x}_{0:T}^{(m)}\}^{M_{1}}_{m=1}$ from the biased MD simulations at the temperature $\lambda_{i}$.
@@
6:        Sample $M_{2}$ paths $\{\bm{x}_{0:T}^{(m)}\}^{M_{2}}_{m=1}$ from the buffer $\mathcal{B}$.
7:        Update parameters of $b_{\theta}$ and $Z_{\theta}$ to minimize the loss $\mathcal{L}_{\text{TB}}$ (Equation 2) using these samples.
@@
To leverage the ability of the TB objective, we use a replay buffer that collects the samples from the tempered version of the generator and reuses them for training. This improves sample efficiency and training stability, both for GFlowNets (Bengio et al., 2021) and reinforcement learning (Mnih et al., 2013). In particular, we collect paths from biased MD simulations with a predefined temperature schedule, starting from a high temperature $\lambda_{\text{start}}$, decreasing the temperature with each rollout, and ending at a temperature $\lambda_{\text{end}}$. This annealing strategy rapidly identifies target metastable states and explores diverse transition paths.
As described in Algorithm 1, the training algorithm of TPS-GFN iterates through four steps: (1) generating paths from the tempered version of the generator, (2) storing generated paths in the replay buffer, (3) sampling paths from the replay buffer, and (4) training the generator by minimizing the TB objective. After training, TPS-GFN directly generates transition paths with biased MD simulation by amortizing inference in the target distribution.
As shown by Malkin et al. (2022b), the TB objective automatically performs gradient variance reduction for variational inference by estimating the normalizing constant $Z$ that acts as a baseline and allows off-policy learning without the need for importance sampling. This is also a slight modification to the VarGrad objective (Richter et al., 2020; Deleu et al., 2024) that was proposed to reduce the variance of the gradient estimator of the evidence lower bound.
We note the similarity of TPS-GFN with the prior work (Holdijk et al., 2024), which minimizes the Kullback-Leibler (KL) divergence between the generator and the target distribution. In comparison, the TB objective yields better training dynamics due to reduced variance in the gradients, and better sample efficiency due to the ability to reuse the collected sample with the replay buffer. Furthermore, TPS-GFN does not suffer from the mode-seeking behavior of KL divergence.
Target distribution relaxation. Learning to imitate the target distribution $p_{\mathcal{A},\mathcal{B}}$ with the TB objective is challenging since its probability is rarely non-zero due to the binary indicator function $\mathbf{1}_{\mathcal{B}}(\bm{r}_{T})$ of the target distribution. To alleviate this issue, we (1) replace the binary indicator function with a piecewise smooth function having positive values for all input paths and (2) allow paths to reach the meta-stable region $\mathcal{B}$ before termination.
First, we relax the target distribution as follows:
@@
where $\tilde{Z}$ is the normalizing constant and $\tilde{\mathbf{1}}_{\mathcal{B}}(\bm{r}_{1:T})$ is the relaxed binary indicator function defined as follows:
@@
$\tilde{\mathbf{1}}_{\mathcal{B}}(\bm{r}_{1:T})=\max_{t\in\{1,\ldots,T\}}\exp\left(-\frac{\lVert D(\bm{r}_{t})-D(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}}{2\sigma^{2}}+h\right),$
where $\lVert\cdot\rVert_{F}$ is the Frobenius norm, $\sigma>0$ controls the degree of relaxation, and $h$ is a scale factor. $D(\bm{r})\in\mathbb{R}^{N\times N}$ is a distance matrix whose $(i,j)$-th element is the Euclidean distance $\lVert\bm{r}_{i}-\bm{r}_{j}\rVert_{2}$ between $i$-th and the $j$-th atoms. Second, we allow the paths $\bm{x}_{1:T}$ to have flexible length, by cutting off the tail of the generated paths of length $T$ at $t_{\ast}=\operatorname{argmin}_{t\in\{1,\ldots,T\}}\lVert D(\bm{r}_{t})-D(\bm{r}_{\mathcal{B}})\rVert_{F}$. This enables finding transition paths of various lengths.
(a) Unbiased MD
@@
Figure 3: 64 sampled paths for each method on the Ramachandran plot of Alanine Dipeptide. White circles indicate meta-stable states, and stars indicate transition states. (3(a)) The paths from unbiased MD simulations that fail to escape the initial meta-stable region. (3(b)) The paths generated by PIPS pass through only one transition state. (3(c)) The paths generated by TPS-GFN pass through both transition states. For clarity, 10 paths are highlighted.
Parameterizing generator with neural bias potential. Similar to Holdijk et al. (2024), we model the generator $p_{\theta}(\bm{x}_{1:T}|\bm{x}_{0})$ as the discretization of the following biased Langevin dynamics:
$\mathrm{d}\bm{r}=\bm{v}\mathrm{d}t,$
$\mathrm{d}\bm{v}=\frac{-\nabla_{\bm{r}}U_{b}(\bm{r})}{\bm{m}}\mathrm{d}t-\gamma\bm{v}\mathrm{d}t+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\mathrm{d}\bm{w},$
where $U_{b}(r)=U(r)+b_{\theta}(\bm{r})$ and $b_{\theta}$ is the neural bias potential that guides the Langevin dynamics towards the meta-stable region $\mathcal{B}$. We discretize the biased Langevin dynamics as follows:
$\bm{r}_{t+1}=\bm{r}_{t}+\bm{v}_{t}\Delta,$
$\bm{v}_{t+1}=(1-\gamma\Delta)\bm{v}_{t}-\frac{-\nabla_{\bm{r}}U_{b}(\bm{r}_{t})}{\bm{m}}\Delta+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\bm{\varepsilon}_{t}\Delta,$
We model the neural bias potential $b_{\theta}(\bm{r})$ or directly force $\mathbf{b}_{\theta}(\bm{r})=-\nabla_{\bm{r}}b_{\theta}(\bm{r})$ for the generator. We use a multi-layered perception (MLP) architecture to parameterize the neural bias potential or force.
4 Experiment
In this section, we evaluate TPS-GFN on three benchmark systems, (1) Alanine Dipeptide, (2) Polyproline Helix, and (3) Chignolin based on realisticity and diversity of generated paths. We compare our method with unbiased MD and path integral stochastic optimal control for path sampling (Holdijk et al., 2024, PIPS) both quantitatively and qualitatively. We run molecular dynamics (MD) simulations, using OpenMM library (Eastman et al., 2023). We report OpenMM and model configurations in Section B.1 and Section B.2, respectively.
4.1 Molecular systems
Alanine Dipeptide. Alanine Dipeptide is a small peptide of two alanine residues connected by a peptide bond. Alanine Dipeptide has the well-known collective variable (CV), the two backbone dihedral angles $(\phi,\psi)$. In our experiments, we sample conformation changes from the meta-stable region $C5$ to $C7_{ax}$ as shown in Figure 2.
Polyproline Helix. Polyproline helix is a protein secondary structure formed by a sequence of proline amino acids (Adzhubei et al., 2013). Polyproline has two meta-stable states: Polyproline I (PP-I) and Polyproline II (PP-II). PP-I is a right-handed helix with all peptide bonds in the cis configuration, while PP-II is a left-handed helix with all peptide bonds in the trans configuration (Moradi et al., 2009). In our experiments, we use Polyproline consisting of 3 proline residues and sample isomerizations from PP-II to PP-I.
Chignolin. Chignolin is a synthetic protein consisting of 10 amino acids and folds into a stable $\beta$-hairpin structure (Honda et al., 2004; Seibert et al., 2005). A network of hydrogen bonds and hydrophobic interactions stabilizes the $\beta$-hairpin structure of chignolin. In our experiments, we sample folding processes from unfolded to folded states.
Figure 4: An isomerization from the meta-stable region PP-II to PP-I of Polyproline generated by TPS-GFN. (Top) 3d views of three states: initial, transition, and final state. The backbone of the Polyproline Helix is highlighted in green. (Middle) The potential energy of states over time. (Bottom) The handedness of states over time. The red line at $y=0$ differentiates between PP-II and PP-I.
@@ 4.2 Quantitative results
We consider three metrics (1) Expected pairwise distance (EPD) (2) target hit percentage (THP) (3) energy transition point (ETP) introduced by Holdijk et al. (2024). EPD measures the similarity between the final conformation and the target conformation by comparing their interatomic distances. THP evaluates how many paths achieve the target meta-stable region in a binary manner. Unlike EPD and THP which only evaluate the final conformations of transition paths, ETP measures the maximum potential energy of a transition path, referring to the ability of a model to find the transition states when crossing the energy barrier.
@@
However, ETP only evaluates the potential energy of the transition state, it can not reflect the realisticity of the overall transition path. To resolve this issue, we propose log-likelihood (LL) which refers to the log-likelihood of conditional density induced by the unbiased MD simulation. Similarly, since EPD and THP only consider the distance between the final conformation of sampled paths and the target conformation, they can not reflect the potential energy of the final conformation which explains its stability and realisticity. To evaluate it, we propose the potential energy of the final point (EFP), and the expected pairwise Coulomb distance (EPCD). EPCD measures both geometric and electrostatic similarity by comparing the Coulomb matrices of the final conformation and the target conformation. The Coulomb matrix captures atomic mass and energy level around the final state better than the distance matrix. More details of metrics are in Section B.3.
@@
Alanine Dipeptide. Since Alanine Dipeptide has the well-known CV, i.e., two backbone dihedral angles $(\phi,\psi)$, we can evaluate THP, ETP, and EFP which requires defining the meta-stable region with the CV. In Table 1, TPS-GFN outperforms the baselines for both previous metrics (EPD, THP, ETP) and proposed metrics (LL, EPCD, EFP) regardless of predicting bias potential or force. The paths generated by our model reach the target meta-stable state better than unbiased MD simulations at 9000K while maintaining the realisticity at 300K. This shows that the TB objective used in our method leads to more realistic and high-quality transition paths than the KL divergence used in PIPS. Thus, our method faithfully amortizes the rejection sampling with unbiased MD simulations at 300K.
@@
Polyproline Helix and Chignolin. Unlike Alanine Dipeptide, the CVs of Polyproline Helix, and Chignolin are less understood. Therefore, we exclude metrics THP, ETP, and EFP which require CVs to define the meta-stable state. In Table 2, our method outperforms baselines in both Polyproline Helix and Chignolin, regardless of predicting bias potential or force. This shows that our model samples more realistic transition paths. We note that predicting bias potential produces better transition paths than predicting bias force since the external force from the bias potential is conservative.
Figure 5: A folding process of Chignolin generated by TPS-GFN. (Top) 3d views of three states: initial, transition, and final state. (Middle) The potential energy over time. (Bottom) The doner-accepter distance of the two key hydrogen bonds, ASP3OD-THR6OG and ASP3N-THR8O over time. To form the hydrogen bonds, the donor-acceptor distance must be lower than the red line at $y=3.5$Å.